OpenStudy (anonymous):

Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours. Use complete sentences

jagr2713 (jagr2713):

First you have to convert all the speed in the same unit.. its entirely upto you to decide.. let us convert speed of SBT train in Japan to m/sec

jagr2713 (jagr2713):

now to convert it we multiply it's speed by 1000/3600 320*1000/1*3600 m/sec.. as 1 km =1000m and 1 hr = 3600 sec

OpenStudy (anonymous):

Are you still stuck on this problem?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

Well as jagr said we need to convert these speeds into m/s. This problem isn't about a specific answer. You are just finding a range for her distance and speed. To to convert km/hr you do this if m is a value of km/hr and n is one for m/s \[(m \frac{ km }{ hr })(\frac{ 1000m }{ 3600s }) = n \frac{ m }{ s }\] So do this for the two speeds and tell me what you get

OpenStudy (anonymous):

Get these values?